Tara pushes a 10.5 kg block of ice 11.0 m across the floor with a force of 235 N. What is the ice block’s velocity the moment she stops pushing it?
1
Expert's answer
2020-11-23T05:29:04-0500

The acceleration given to the block is (according to Newton's second law):


"a = \dfrac{F}{m}"

where "m = 10.5kg" is the mass of the block and "F = 235N" is the exerted force.

Assuming block started from rest, the distance travelled by it in time "t" is:


"x = \dfrac{at^2}{2}"

As far as "x = 11m", this time will be:


"t = \sqrt{\dfrac{2x}{a}} = \sqrt{\dfrac{2xm}{F}}"

The speed gained in this time is:


"v = at = \dfrac{F}{m}\sqrt{\dfrac{2xm}{F}} = \sqrt{\dfrac{2xF}{m}}\\
v = \sqrt{\dfrac{2\cdot 11\cdot 10.5}{235}} \approx 0.99 m/s"

Answer. 0.99 m/s.